OpenStudy (love_to_love_you):

Please help, will fan and medal! Solve the system using an augmented matrix. Show all work. -x-3y=-17 2x-6y=-26

OpenStudy (love_to_love_you):

@dan815

OpenStudy (love_to_love_you):

I really really need help.

OpenStudy (love_to_love_you):

@aum

OpenStudy (love_to_love_you):

@Ilovecake

OpenStudy (love_to_love_you):

@campbell_st

OpenStudy (campbell_st):

well doesn't that mean \[\left[\begin{matrix}-1 & -3 \\ 2& -6\end{matrix}\right] = \left[\begin{matrix}-17 \\ -26\end{matrix}\right]\] then its the 1st row times 2 added to the 2nd row which gives \[\left[\begin{matrix}-1 & -3\\ 0 & -12\end{matrix}\right] =\left[\begin{matrix}-17 \\ -60\end{matrix}\right]\] divide row 2 by -12 \[\left[\begin{matrix}-1 & 3 \\ 0 & 1\end{matrix}\right] =\left[\begin{matrix}-17\\ 5\end{matrix}\right]\] This is just a guess, not sure if you need \[\left[\begin{matrix}x \\ y\end{matrix}\right]\] there as well. hope it helps

OpenStudy (love_to_love_you):

thank youuuuu